# ProReason: Multi-Modal Proactive Reasoning with Decoupled Eyesight and Wisdom

### v2

@@ 3.3 Advantages of ProReason
In ProReason, the complete multi-modal reasoning process is decomposed into visual perception and textual reasoning stages, each executed by separate agents. These agents are then effectively organized through a designated pipeline, significantly enhancing the ability of LVLMs to tackle visual reasoning tasks.
@@
On the other hand, this decomposition-before-integration approach not only circumvents the inherent differences between modalities but also allows visually irrelevant sub-agents to be performed by text-only LLMs, thereby facilitating LLM-assisted multi-modal reasoning. Consequently, the extensively verified powerful reasoning capabilities of LLMs can be seamlessly synergized to power the multi-modal reasoning process for performance enhancement.
@@
With the close collaboration of the sub-agents in Proactive Visual Perception step, ProReason can proactively (i.e., question-orientedly and reasoning-involvedly) extract the necessary and sufficient visual details from images, effectively avoiding information omission or redundancy. Meanwhile, the Memory component retains only the image captions from the Vision Expert and the reasoning results from the Reasoning Expert, providing a compact textual descriptions. This contributes to minimize the interference of irrelevant information on the subsequent Summarizer.
@@ 4.1 General Setup
To comprehensively validate the performance of our framework, we conduct experiments across four benchmarks: Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a)  [^3] Due to our emphasis on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning. To facilitate the comparison across different benchmarks, the results for MME benchmark are calculated by the percentage of correct answers out of the total answers. For the results using the official MME calculation method, please refer to Table 16, Table 16, Table 18, Table 16 and Table 18. , Massive Multi-discipline Multi-modal Understanding and Reasoning (MMMU) (Yue et al., 2023), MathVista (Wang et al., 2024), and HallusionBench (Liu et al., 2023a). All of them require visual reasoning capabilities to complete the tasks correctly, and are introduced briefly in Sec. A.2.
@@
We employ GPT-4o-mini, Llama3-LLaVA-NeXT-8B [^4] The model is deployed using CUDA https://developer.nvidia.com/cuda-toolkit on a NVIDIA A100 graphics cards. and Qwen2.5-72B- Instruct (Team, 2024) as our base models, due to their excellent performance, accessibility, and representativeness of model sources. As one of the most performant LVLMs, GPT-4o-mini demonstrates significant advancements in visual reasoning capabilities, and provides cheap and fast API. In contrast, Llama3-LLaVA-NeXT-8B is a fully open-sourced LVLM developed by the LLaVA team based on the Llama-3-8B LLM (Dubey et al., 2024) and CLIP vision encoder (Radford et al., 2021). Qwen2.5-72B-Instruct is a robust LLM designed to deliver high-quality instruction following and handle complex tasks. It will be utilized for achieving LLM-assisted multi-modal reasoning.
@@
Besides the most basic method where models are instructed to answer questions directly, we compare ProReason with two categories of peer methods. To determine the benefits of proactive information extraction in ProReason, we first consider two SOTA passive visual reasoning methods, VDGD (Ghosh et al., 2024) and CCoT (Mitra et al., 2024). Additionally, we choose two multi-step reasoning frameworks of LLMs, Chain of Thought (CoT) (Wei et al., 2022) and ReAct (Yao et al., 2022), to demonstrate the effects of directly migrating LLM solutions to LVLMs.
  2. VDGD. With image caption prefixed to text instruction, VDGD selects tokens that deviate the least from the description using a formula based on Kullback-Leibler divergence.
  5. ReAct. ReAct is an LLM-specific agent framework, which performs tasks by alternating between reasoning and execution behaviors. To extend it to multi-modal domain, we use two LVLMs to perform both steps, and rename them as the Vision and Reasoning Experts, respectively. This aligns with our notions for easy understanding, and is shown in Figure 1.
@@
The prompt templates for all methods are shown in Figures 9, 10 and 12. Specifically for ProReason, to prevent infinite loops, if the Dispatcher selects the Vision Expert or Reasoning Expert to obtain information up to 5 consecutive times, and the Referee still determines that the existing knowledge in Memory remains insufficient to resolve the question, the Memory will be cleared to restart the information acquisition process. If the system fail to break the loop after 5 attempts, the proactive visual perception phase will be immediately terminated. Subsequently, the Summarizer will generate the final answer based on the available information in Memory from the last attempted iteration. This setup, refined through multiple trials, is the most effective. Additionally, since we cannot obtain the tokens output by GPT-4o-mini, we omit the step of selecting the token with the smallest deviation from the image description when implementing VDGD for GPT-4o-mini.
Table 1: Performance of multiple approaches with two base models across four visual reasoning benchmarks. “Hallu.” is the abbreviation of HallusionBench. Based on the performance of the direct method, red and blue signify the improvement and degradation, respectively.
@@
MME MMMU MathVista Hallu.
Llama3-LLaVA- NeXT-8B Direct 61.5 41.8 37.1 45.8
VDGD 68.8 (+7.3) 42.3 (+0.5) 36.1 (-1.0) 44.2 (-1.6)
CCoT 68.9 (+7.4) 40.5 (-1.3) 36.8 (-0.3) 37.4 (-8.4)
CoT 58.8 (-2.7) 41.5 (-0.3) 35.9 (-1.2) 43.1 (-2.7)
ReAct 68.5 (+7.0) 46.7 (+4.9) 31.7 (-5.4) 43.6 (-2.2)
ProReason 71.5 (+10.0) 50.5 (+8.7) 38.8 (+1.7) 50.9 (+5.1)
Average 66.30 43.9 36.06 44.16
GPT-4o-mini Direct 79.2 48.4 53.0 56.0
VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6)
CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7)
CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3)
ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9)
ProReason 91.9 (+12.7) 61.6 (+13.2) 54.9 (+1.9) 59.9 (+3.9)
Average 84.63 54.82 52.63 55.4
Table 2: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks. The red and blue texts indicate the improvements and reductions relative to the case of using the same method without the assistance of text-only LLMs, respectively.
Vision Expert Textual Sub-Agents Method Dataset
MME MMMU MathVista Hallu.
Llama3- LLaVA- NeXT-8B Llama3-LLaVA- NeXT-8B Direct 61.5 41.8 37.1 45.8
ReAct 68.5 46.7 31.7 43.6
ProReason 71.5 50.5 38.8 50.9
Qwen2.5-72B- Instruct ReAct 71.0 (+2.5) 50.4 (+3.7) 34.6 (+2.9) 40.4 (-3.2)
ProReason 81.3 (+9.8) 56.8 (+6.3) 48.8 (+10.0) 52.3 (+1.4)
GPT-4o-mini ReAct 73.6 (+5.1) 48.4 (+1.7) 36.2 (+4.5) 46.7 (+3.1)
ProReason 84.7 (+13.2) 54.5 (+4.0) 41.7 (+2.9) 53.1 (+2.2)
@@ 4.2 Main Results
ProReason exhibits significant and consistent performance enhancement over baselines across all the benchmarks. As listed in Table 1, despite better performance than the direct method on MME dataset, VDGD and CCoT fail to demonstrate consistent improvements on the other datasets. In contrast, ProReason consistently outperforms all the other baselines for both Llama3-LLaVA-NeXT-8B and GPT-4o-mini model across all benchmarks, with a peak improvement as 13.2%, demonstrating the superiority and task robustness of ProReason.
Proactive information acquisition surpasses SOTA passive methods, especially in complex visual reasoning tasks. Specifically, compared to MME, MathVista and HallusionBench present higher image complexity and question difficulty, and thus require stronger visual understanding and textual reasoning capabilities. This leads to performance degradation of passive methods (i.e., VDGD and CCoT), highlighting their limited applicability to complex visual reasoning tasks. In contrast, ProReason achieves notable performance improvements, up to 5.1%, by proactively acquiring visual information from images rather than generating question-agnostic captions. This aligns with out previous obervations in Sec. 2.2 that passive methods introduce substantial information redundancy or omission, misleading subsequent reasoning processes.
@@
Decoupling the visual perception and textual reasoning capabilities of an LVLM outperforms their simultaneous inherent usage. As listed in Table 1, when both capabilities are utilized concurrently, CoT consistently degrades performance compared to the “Direct” method across all benchmarks with Llama3-LLaVA-NeXT-8B model, consistent with the findings in Sec. 2.1. In contrast, despite the same models, ProReason alternates between visual information acquisition and textual reasoning processes, allowing to leverage each capability more effectively. This enables ProReason to consistently outperform CoT with both Llama3-LLaVA-NeXT-8B and GPT-4o-mini across all benchmarks, demonstrating the effectiveness of capability decoupling.
ProReason outperforms ReAct with even less token consumption. Given the similarity in multi-step reasoning, we compare ProReason and ReAct in terms of performance and cost. Specifically, as an LLM-specific multi-step reasoning framework, ReAct only outperforms the “Direct” method on MME and MMMU, but underperforms on MathVista and HallusionBench, showing inferior performance compared to the consistent improvements of ProReason. Furthermore, we compare their average token consumption of GPT-4o-mini. As shown in Table 13 of Appendix A, ProReason consumes significantly fewer tokens than ReAct on both MME and MathVista tasks, indicating its higher token efficiency and the importance of a compact Memory in reducing token usage. Coupled with better performance, this suggests the superiority of ProReason over LLM-specific ReAct framework.
@@
Text-only LLMs can be effectively integrated into ProReason for dramatically enhanced performance. As mentioned in Sec. 3.3, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs. To demonstrate the utility of this advantage, we fix the Vision Expert as Llama3-LLaVA-NeXT-8B, and replace other agents with text-only LLMs. As listed in Table 2, with the assistance of either Qwen2.5-72B-Instruct or GPT-4o-mini, the Llama3-LLaVA-NeXT-8B Vision Expert receives remarkable performance boost across all benchmarks, particularly by 15% on MMMU and 11.7% on MathVista, compared to directly providing answers. In contrast, ReAct gains a much smaller improvement. This highlights the unique advantage of ProReason in leveraging existing text-only LLMs for enhanced performance. Notably, this advantage may open new avenues for continuously pushing the performance limits of LVLMs with the assistance of existing powerful LLMs.
4.3 Relative Importance of Sub-agents
To assess the importance of each sub-agent within the ProReason framework for visual reasoning tasks, we design five scenarios where Llama3-LLaVA-NeXT-8B acts as Dispatcher, Vision Expert, Reasoning Expert, Referee, or Summarizer, respectively, while the other sub-agents are powered by GPT-4o-mini. Given that Llama3-LaVA-NeXT-8B exhibits weaker visual understanding and textual reasoning capabilities than GPT-4o-mini, the more significant the performance drop incurred by replacing a sub-agent with Llama3-LaVA-NeXT-8B is, the more important that sub-agent is. Here we primarily consider the MME and MMMU benchmarks due to their comprehensive question coverage. The experimental results are presented in Table 13.
Relatively, Dispatcher and Reasoning Expert are the least essential sub-agents. Specifically, despite a decline, these two sub-agents exhibit significantly less performance degradation than other sub-agents. This can be attributed to the easier task of the Dispatcher, which requires minimal textual reasoning capabilities, and the infrequent calls of the Reasoning Expert, which is only activated when additional information needs to be inferred—a situation that is rare in current benchmarks. Besides, both sub-agents operate within the acquisition loop, allowing for greater error tolerance. Even if some error occurs, subsequent iterations can compensate for the missing information.
In summary, each sub-agent contributes to the performance of ProReason, underscoring their necessity. Relatively, the Summarizer and Referee are the most critical sub-agents, while the Dispatcher and Reasoning Expert have the least impact.
4.4 The Critical Implication of Decoupling
  1. Merge the Vision Expert and Reasoning Expert into a single sub-agent to examine the necessity of modality decoupling during sub-task execution.
  2. Integrate the Dispatcher, Vision Expert, Reasoning Expert, and Referee as a unified agent to verify the essentiality of the process design for the Proactive Visual Perception stage.
Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks. As shown in Table 13, merging Vision and Reasoning Experts results in a 3% performance drop on the MMMU benchmark, while combining Proactive Visual Perception with Textual Reasoning leads to a more significant 4.8% reduction (56.8 vs. CoT’s 58.5). Although agent merging also causes performance declines on MME, the merged versions still outperform CoT. Notably, given MMMU’s substantially higher complexity compared to MME, these findings reveal that the decoupling of visual-textual processing fundamentally drives its performance gains in complex scenarios.

### v3

@@ 3.3 Advantages of ProReason
In ProReason, the multi-modal reasoning process is decomposed into visual perception and textual reasoning stages, each executed by separate agents. These agents are then effectively organized through a designated pipeline, significantly enhancing the ability of LVLMs to tackle visual reasoning tasks.
@@
Decoupled vision-reasoning eliminates the necessity for LVLMs to manage vision-irrelevant sub-agents, allowing the seamless integration of existing LLMs with established strong reasoning abilities, thereby endowing ProReason with superior visual reasoning performance.
Downstream task model enhancement.
By generating high-quality data, ProReason effectively distills its superior capabilities into downstream task models, enabling them to demonstrate exceptional performance in complex visual reasoning tasks.
@@
Through sub-agent collaboration in Proactive Visual Perception, ProReason extracts essential visual details via question-oriented manner, preventing information omission or redundancy. Meanwhile, the Memory component retains only Vision Expert’s observations and Insight Expert’s conclusions, creating compact representation that minimizes irrelevant information. These advantages effectively reduce the overhead when ProReason functions as a reasoning framework.
@@ 4.1 General Setup
To comprehensively validate the performance of our framework, we conduct experiments across four benchmarks: Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023)  [^2] Due to our emphasis on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning. To facilitate the comparison across different benchmarks, the results for MME benchmark are calculated by the percentage of correct answers out of the total answers. , Massive Multi-discipline Multi-modal Understanding and Reasoning (MMMU) (Yue et al., 2023), MathVista (Wang et al., 2024), and HallusionBench (Liu et al., 2023a). All of them require visual reasoning capabilities to complete the tasks correctly, and are introduced briefly in Section A.2.
@@
We employ GPT-4o-mini, Llama3-LLaVA-NeXT-8B and Qwen2.5-VL-7B- Instruct (Bai et al., 2025) [^3] The model is deployed using CUDA https://developer.nvidia.com/cuda-toolkit on a NVIDIA A100 graphics cards. to drive ProReason, and Qwen2.5-72B-Instruct (Yang et al., 2024) and Qwen3-32B (Team, 2025) are selected to drive the text-only sub-agents, enabling LLM-assisted visual reasoning. In addition Qwen2.5-VL-7B-Instruct is also utilized as a downstream task model, which is trained with the data generated by ProReason. These models are chosen for their leading performance and popularity.
@@
In addition to the basic method where models are instructed to answer questions directly, we compare ProReason with the following peer methods. First, to explore the effectiveness of directly migrating LLM solutions to LVLMs, we select two multi-step reasoning frameworks from LLMs: Chain of Thought (CoT) (Wei et al., 2022) and ReAct (Yao et al., 2022). Second, to evaluate the advantages of proactive information extraction in ProReason, we examine two passive visual reasoning frameworks, VDGD (Ghosh et al., 2024) and CCoT (Mitra et al., 2024), which assist LVLM reasoning by extracting image information into text. Additionally, we consider R1-Onevision-7B (Yang et al., 2025), an LVLM based on Qwen2.5-VL-7B-Instruct with deep thinking capabilities similar to Deekseek-R1 (Guo et al., 2025) and OpenAI o1444https://openai.com/index/introducing-openai-o1-preview/. We compare it with ProReason-Q3, which also exhibits deep thinking abilities. Further details on the baselines and implementation are provided in Section A.3.
@@
The prompt templates for all methods are shown in Figures 9, 10 and 12. Specifically for ProReason, to prevent infinite loops, if the Dispatcher selects the Vision Expert or Insight Expert to obtain information up to 5 consecutive times, and the Referee still determines that the existing knowledge in Memory remains insufficient to resolve the question, the Memory will be cleared to restart the information acquisition process. If the system fail to break the loop after 5 attempts, the proactive visual perception phase will be immediately terminated. Subsequently, the Summarizer will generate the final answer based on the available information in Memory from the last attempted iteration. This setup, refined through multiple trials, is the most effective.
Table 1: Performance of multiple approaches with three base models across four visual reasoning benchmarks. “Hallu.” is the abbreviation of HallusionBench. The abbreviations 4o-mini, Qwen72B, and Qwen3 refer to GPT-4o-mini, Qwen2.5-72B-Instruct, and Qwen3-32B, respectively. "Assisted" stands for LLM-assisted reasoning, which involves replacing the textual sub-agents within frameworks with corresponding LLMs, as detailed in Table 14. Based on the performance of the direct method, red and blue signify the improvement and degradation, respectively.
@@
MME MMMU MathVista Hallu. Average
GPT-4o-mini Direct 79.2 48.4 53.0 56.0 59.2
VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6) 59.3 (+0.1)
CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7) 61.3 (+2.1)
CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3) 64.1 (+4.9)
ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9) 60.6 (+1.4)
ProReason 91.9 (+12.7) 61.6 (+13.2) 54.9 (+1.9) 59.9 (+3.9) 67.1 (+7.9)
Llama3-LLaVA-NeXT-8B Direct 61.5 41.8 37.1 45.8 46.6
VDGD 68.8 (+7.3) 42.3 (+0.5) 36.1 (-1.0) 44.2 (-1.6) 47.8 (+1.2)
CCoT 68.9 (+7.4) 40.5 (-1.3) 36.8 (-0.3) 37.4 (-8.4) 45.9 (-0.7)
CoT 58.8 (-2.7) 41.5 (-0.3) 35.9 (-1.2) 43.1 (-2.7) 44.8 (-1.8)
ReAct 68.5 (+7.0) 46.7 (+4.9) 31.7 (-5.4) 43.6 (-2.2) 47.6 (+1.0)
+ 4o-mini Assisted 73.6 (+12.1) 48.4 (+6.6) 36.2 (-0.9) 46.7 (+0.9) 51.2 (+4.6)
+ Qwen72B Assisted 71.0 (+9.5) 50.4 (+8.6) 34.6 (-2.5) 40.4 (-5.4) 49.1 (+2.5)
ProReason 71.5 (+10.0) 50.5 (+8.7) 38.8 (+1.7) 50.9 (+5.1) 52.9 (+6.3)
+ 4o-mini Assisted 84.7 (+23.2) 54.5 (+12.7) 41.7 (+4.6) 53.1 (+7.3) 58.5 (+11.9)
+ Qwen72B Assisted 81.3 (+19.8) 56.8 (+15.0) 48.8 (+11.7) 52.3 (+6.5) 59.8 (+13.2)
Qwen2.5-VL-7B-Instruct Direct 74.2 51.8 63.3 53.8 60.8
VDGD 74.6 (+0.4) 52.3 (+0.5) 62.1 (-1.2) 53.9 (+0.1) 60.7 (-0.1)
CCoT 82.7 (+8.5) 52.2 (+0.4) 61.7 (-1.6) 55.8 (+2.0) 63.1 (+2.3)
CoT 72.3 (-1.9) 53.6 (+1.8) 63.7 (+0.4) 55.9 (+2.1) 61.4 (+0.6)
ReAct 81.9 (+7.7) 51.6 (-0.2) 60.4 (-2.9) 52.9 (-0.9) 61.7 (+0.9)
+ Qwen72B Assisted 83.1 (+8.9) 52.5 (+0.7) 62.2 (-1.1) 54.5 (+0.7) 63.1 (+2.3)
ProReason 83.8 (+9.6) 57.0 (+5.2) 63.2 (-0.1) 56.4 (+2.6) 64.6 (+3.8)
+ Qwen72B Assisted 92.7 (+18.5) 64.6 (+12.8) 64.0 (+0.7) 60.6 (+6.8) 70.5 (+9.7)
+ Qwen3 Assisted 90.7 (+16.5) 66.2 (+14.4) 67.2 (+3.9) 59.8 (+6.0) 71.0 (+10.2)
@@ 4.2 Main Results
ProReason exhibits significant and consistent performance enhancement over baselines across all the benchmarks. As listed in Table 1, despite better performance than the direct method on MME dataset, VDGD and CCoT fail to demonstrate consistent improvements on the other datasets. In contrast, ProReason consistently surpasses all other baselines across all benchmarks for every base model, enhancing the average performance of GPT-4o-mini by 7.9%, demonstrating the superiority and task robustness of ProReason. Section A.5.7 and Table 13 evaluate ProReason on additional benchmarks, including reasoning and VQA tasks, further proving its strong performance and effectiveness across various multi-modal tasks.
@@
Decoupling the visual perception and textual reasoning capabilities of an LVLM outperforms their simultaneous inherent usage. Table 1 illustrates that CoT, utilizing both capabilities simultaneously, does not exhibit consistent performance enhancements over the "Direct" method when applied to the Llama3-LLaVA-NeXT-8B and Qwen2.5-VL-7B-Instruct models. In contrast, despite the same models, ProReason alternates between visual information acquisition and textual reasoning processes, allowing to leverage each capability more effectively. This enables ProReason to consistently outperform CoT with both Llama3-LLaVA-NeXT-8B and GPT-4o-mini across all benchmarks, validating the effectiveness of capability decoupling.
Proactive information acquisition surpasses peer passive methods, especially in complex visual reasoning tasks. Specifically, compared to MME, MathVista and HallusionBench present higher image complexity and question difficulty, and thus require stronger visual understanding and textual reasoning capabilities. This leads to performance degradation of passive methods (i.e., VDGD and CCoT), highlighting their limited applicability to complex visual reasoning tasks. In contrast, ProReason achieves notable performance improvements, up to 5.1%, by proactively acquiring visual information from images rather than generating question-agnostic captions. This aligns with out previous obervations in Section 2 that passive methods introduce substantial information redundancy or omission, misleading subsequent reasoning processes.
@@
Text-only LLMs can be effectively integrated into ProReason for dramatically enhanced performance. As mentioned in Section 3.3, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs. To demonstrate the utility of this advantage, we fix the Vision Expert and replace the textual sub-agents in ProReason with text-only LLMs, according to the configuration in Table 14. As listed in Table 1, with the assistance of powerful existing LLMs, the Llama3-LLaVA-NeXT-8B Vision Expert receives remarkable performance boost across all benchmarks, particularly by 15% on MMMU and 11.7% on MathVista, compared to directly providing answers. In addition, by configuring the Summarizer to Qwen3-32B, ProReason acquires the same kind of deep reasoning capability as Deekseek-R1 and OpenAI o1, resulting in a 10.2% average performance improvement for the Qwen2.5-VL-7B-Instruct Vision Expert. In contrast, ReAct gains a much smaller improvement. This highlights the unique advantage of ProReason in leveraging existing text-only LLMs for enhanced performance. Notably, this advantage may open new avenues for continuously pushing the performance limits of LVLMs with the assistance of existing powerful LLMs.
Table 2: Performance of different models across two visual benchmarks. Based on the performance of the base model(i.e., Qwen2.5-VL-7B-Instruct), red and blue signify the improvement and degradation, respectively.
Dataset Model
Qwen2.5-VL-7B-Instruct R1-Onevision-7B ProReason-VL ProReason-Q3
MME 72.3 88.1 (+15.8) 91.2 (+18.9) 90.8 (+18.5)
MMMU(val.) 53.6 53.1 (-0.5) 65.4 (+11.8) 67.4 (+13.8)
Average 63.0 70.6 (+7.6) 78.3 (+15.3) 79.1 (+16.1)
Table 3: Average token and time consumption of multiple approaches with GPT-4o-mini model on the MME and MathVista benchmarks.
Dataset Method GPT-4o-mini
Direct VDGD CCoT CoT ReAct ProReason
MME Input 393.9 1020.9 1024.3 403.9 1645.0 1286.8
Output 5.9 155.6 254.3 103.4 197.0 327.2
Time(s) 6.1 12.3 12.8 7.3 18.9 18.4
MathVista Input 368.4 955.5 961.2 375.4 3092.8 2238.6
Output 51.8 263.3 307.3 479.7 845.1 788.6
Time(s) 4.2 11.5 12.3 12.0 28.8 24.8
4.3 Downstream Task Model Enhancement
High-quality visual reasoning data produced by ProReason significantly improves the performance of downstream task models. We select the test set that has no answer in the MMMU dataset and gather reasoning processes paired with corresponding answers on this set, generated by two configurations detailed in Table 14: ProReason + Qwen72B Assisted and ProReason + Qwen3 Assisted. After eliminating samples with inconsistent answers between these two configurations, we obtain two filtered datasets each containing 5,980 entries. These datasets are then used to fine-tune Qwen2.5-VL-7B-Instruct (Section A.4), yielding two models: ProReason-VL and ProReason-Q3. Notably, ProReason-Q3 inherits the deep reasoning capabilities of Qwen3-32B, which incorporates a reasoning process containing a <think>... </think> part.
Table 2 illustrates that while R1-Onevision-7B features deep reasoning capabilities and shows performance enhancements on MME, it fails to demonstrate improvement on the MMMU validation set. In contrast, both ProReason-VL and ProReason-Q3 achieve performance gains exceeding 10% on the MMMU validation set and also exhibit significant improvements on MME. This indicates that ProReason effectively transfers the exceptional reasoning capabilities of LLM-integrated decoupled systems to downstream task models, highlighting the potential for leveraging the robust abilities of existing LLMs to continuously enhance the performance of LVLMs.
4.4 Efficiency and Complexity Analysis
ProReason improves both efficiency and performance compared to the baseline method. Table 3 presents the evaluation of average token consumption and time expenditure of different methods on the MME and MathVista datasets. Notably, ProReason requires significantly fewer tokens than ReAct, while achieving superior performance over ReAct as analyzed in Section 4.2. Moreover, compared to the visual reasoning frameworks VDGD and CCoT, which also involve multiple image inputs, ProReason’s token consumption on MME is only about 20% higher, yet it achieves a 12.7% performance improvement. Regarding time efficiency, ProReason achieves 11.1% higher performance than VDGD and CCoT on MME with comparable time consumption (18.4s/sample vs. VDGD’s 12.3s and CCoT’s 12.8s), while ReAct is slower than ProReason across datasets, highlighting ProReason’s efficiency advantage. Furthermore, as discussed in Section 4.3, ProReason’s robust visual reasoning capabilities can be transferred to downstream task models, further ensuring the performance and efficiency of our approach.
ProReason achieves a balance between performance and efficiency through its Memory component and adaptive mechanisms. The Memory component enables ProReason to keep compact information representations, avoiding lengthy reasoning traces like ReAct, reducing token use and boosting efficiency. Additionally, ProReason dynamically adjusts Proactive Visual Perception iterations based on question difficulty, minimizing overhead for simple tasks while enhancing complex problem-solving. As shown in Table 12, MathVista’s greater challenge prompts more iterations compared to MME, resulting in higher token usage and longer reasoning times. This aligns with the difficulty levels of the datasets, demonstrating the adaptive nature of ProReason.
4.5 Ablation Study and Further Analysis
Relative Importance of Sub-agents.
To evaluate the importance of each sub-agent in ProReason, in Section A.5.1, we replace Dispatcher, Vision Expert, Insight Expert, Referee, and Summarizer individually with the less capable Llama3-LLaVA-NeXT-8B(which demonstrates weaker visual understanding and textual reasoning capabilities), while keeping the other sub-agents as GPT-4o-mini. The performance degradation on the MME and MMMU benchmarks is then used to measure the significance of each sub-agent. Results indicate that the Summarizer is the most critical sub-agent, closely followed by Referee.
Which One is More Crucial: Visual Understanding or Textual Reasoning
In Section A.5.2, we perform comparative experiments by substituting the vision expert and text sub-agents in the ProReason with Llama3-LLaVA-NeXT-8B and GPT-4o-mini respectively. Our findings indicate that while both visual understanding and textual reasoning capabilities are essential for multimodal tasks, textual reasoning ability holds greater significance in visual reasoning tasks. This result is consistent with our earlier analysis in Section A.5.1, which identifies the Summarizer and Referee as the most critical sub-agents.
The Critical Implication of Decoupling.
In Section A.5.3, to validate the necessity of decoupling visual perception and textual reasoning in ProReason, we systematically merge sub-agents through three configurations while preserving identical prompts and procedures, with observed performance degradation quantitatively demonstrating the critical role of decomposed processing in enhancing capabilities. The experimental results indicate that Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks
Reasoning Process Evaluation of ProReason.
In Section A.5.4, we evaluate the responses generated by ProReason using LLMs. The analysis reveals that, compared to CoT, ProReason produces more relevant answers with reduced redundancy and deficiency, consistent with its enhanced performance.
Referee’s Dispel of Hallucinations.
In Section A.5.5, adhering to the settings in Section 4.1, we assess ProReason (powered by GPT-4o-mini) on MMMU and HallusionBench with different attempt allowances (1/3/5). When attempts are unsuccessful, systematic Memory clearance is triggered (determined by the Referee’s lack of sufficient information). As the number of attempts increases, the Referee has more opportunities for information filtering. The observed performance variations illustrate the crucial influence of the Referee’s decision - making and filtering efficiency on the system’s capabilities. Experimental results indicate that the Referee module effectively filters hallucinated information to improve the visual comprehension capabilities of our framework.
Frequency of selection of various experts.
In Section A.5.6, we assess how often the Dispatcher selects the Vision Expert or Insight Expert across both MME and MMMU benchmarks. The experimental results demonstrate that ProReason adaptively adjusts the frequencies of expert selection, leading to consistent performance improvements.